# Adaptive Online Bayesian Estimation of Frequency Distributions with Local Differential Privacy

### v1

@@ 8 Conclusion
In this paper, we proposed a novel adaptive framework, AdOBEst-LDP, for online estimation of the frequency distribution of categorical data under the $\epsilon$-LDP constraint. AdOBEst-LDP, run with RRRR for randomization, encompasses both the privatization/perturbation of the given data and the accurate estimation of population parameters from privatized data, in a dynamic way, using Bayesian methodology. Our privatization mechanism (RRRR) is distinguished from the baseline approach (SRR) in a way that it operates on a smaller subset of the categories rather than using all categories which is often suboptimal in terms of accuracy. Finding the optimal subset for this task in a static way is a computationally inefficient approach whose complexity is in the order of $2^{K}$. For that reason, we employed an adaptive approach to dynamically update the subsets at each iteration, given the current sample of the population parameter and the privacy parameter. Optimal selection of these subsets was guided by various alternative utility functions that we used throughout the paper, and in each of them, their ‘utility’ was based on optimizing a different information metric to select the most informative subset among all possible combinations, in a computationally tractable way.
For the posterior sampling of $\theta$ at each iteration, we employed an efficient SGLD-based sampling scheme on a constrained region, namely the $K$-dimensional probability simplex. We distinguished this scheme from Gibbs sampling which uses all of the historical data, which is not scalable to large datasets.
@@
In the numerical experiments, we demonstrated that AdOBEst-LDP can estimate the population distribution more accurately than the non-adaptive approach under experimental settings with various privacy levels $\epsilon$ and degrees of evenness among the components of $\theta^{\ast}$. While the performance of AdOBEst-LDP is generally robust for all the utility functions considered in the paper, the utility function based on the probability of honest response can be preferred due to its much lower computational complexity than the other utility functions. Our experiments also showed that the accuracy of the adaptive approach is comparable to that of the semi-adaptive approach. However, the semi-adaptive approach requires determining an optimal parameter ($\alpha$) in a continuous space, so the adaptive approach is practically more useful in real-life applications.
@@
Via theoretical analysis, we showed that, regardless of whether the posterior sampling is conducted exactly or approximately, the target posterior distribution of AdOBEst-LDP converges to the true population parameter $\theta^{\ast}$. We also showed that, under exact posterior sampling, the best subset given utility function is selected with probability $1$ in the long run.
@@
Note that the observations $\{Y_{t}\}_{t\geq 1}$ generated by AdOBEst-LDP are dependent. Hence, the theoretical work in Section 6 can also be regarded as a contribution to the literature on the convergence of posterior distributions under dependent observations. Moreover, we have already pointed out a form of analogy between AdOBEst-LDP and Thompson sampling (Russo et al., , 2018). Observe that posterior sampling is present in both methods and the subset selection step can be likened to the action step of a reinforcement learning scheme. In that sense, we also believe that the theoretical results can inspire studies on the convergence of dynamic reinforcement learning algorithms, particularly Thompson sampling.
Appendix A Appendix
@@ A.1 Proofs for LDP of RRRR
for all possible $x,x^{\prime},y\in\{1,\ldots,K\}$. When $S=\emptyset$, the proof is trivial; we focus on the non-trivial case $S\neq\emptyset$. For the non-trivial case, the transition probability $g_{S,\epsilon}(y|x)$ requires checking the ratio in (22) in $10$ different cases for $x,x^{\prime},y$ concerning their interrelation.
@@
    Since $\epsilon_{1}<\epsilon$, (22) holds.
@@
    hence we proved the right-hand side inequality. For the left-hand side, we have
@@
    Since $\epsilon_{1}<\epsilon$, (22) holds.
@@
    So (22), trivially holds.
We conclude the proof by noting that any other case left out is symmetric in $(x,x^{\prime})$ to one of the covered cases and therefore does not need to be checked separately. ∎
@@ A.2 Proofs about utility functions
$F(\theta;S,\epsilon)=\sum_{y=1}^{K}\frac{a_{y}^{\top}}{h_{S,\epsilon}(y|\theta)}\frac{a_{y}}{h_{S,\epsilon}(y|\theta)}h_{S,\epsilon}(y|\theta)=\sum_{y=1}^{K}a_{y}^{\top}\frac{1}{h_{S,\epsilon}(y|\theta)}a_{y}=A_{S,\epsilon}^{\top}D_{\theta}A_{S,\epsilon}$
@@
$G_{S,\epsilon}$ is invertible for all $S\subset\{1,\ldots,K\}$ and $\epsilon>0$.
@@
The case $k=0$ is trivial since then $G_{S,\epsilon}=d_{1}I_{K}+d_{2}1_{K}1_{K}^{\top}$ is invertible. Hence, we focus on the case $0<k<K$. For this case, firstly note that the matrices on the diagonal are invertible. So, by Weinstein–Aronszajn identity, for $G_{S,\epsilon}$ to be invertible, it suffices to show that the matrix
@@
Inside the square brackets is a scalar, therefore $M$ in question is the sum of an identity matrix and a rank-$1$ matrix, which is invertible. Hence, $G_{S,\epsilon}$ is invertible. ∎
@@
where $c^{G}_{i}$ is the $i$’th column of $G_{S,\epsilon}$ for $i=1,\ldots,K$. Observe that $c^{A}_{i}$, $i=1,\ldots,K-1$ are linearly independent, since any linear combination of those columns is in the form of
@@
Since the columns of $G_{S,\epsilon}$ are linearly independent, the linear combination above becomes 00 only if $a_{1}=\ldots=a_{K-1}=0$. This shows that the columns of $A_{S,\epsilon}$ are also linearly independent, thus we conclude that $A_{S,\epsilon}$ has rank $K-1$. Finally, since $D_{\theta}$ is diagonal with positive diagonal entries, $A_{S,\epsilon}^{\top}D_{\theta}A_{S,\epsilon}=A_{S,\epsilon}^{\top}D_{\theta}^{1/2}D_{\theta}^{1/2}A_{S,\epsilon}$ is positive definite, hence invertible. ∎
@@
The global maximization of $U_{6}$ over all $S\subset\{1,\ldots,K\}$ can be decomposed as
$\max_{S\subset\{1,\ldots,K\}}U_{6}(\theta,S,\epsilon)=\max_{k\in\{0,\ldots,K-1\}}\left\{\max_{S\subset\{1,\ldots,K\}:|S|=k}U_{6}(\theta,S,\epsilon)\right\}.$ (28)
@@ A.4.1 Preliminary results
Given $\epsilon\geq 0$, there exists constants $0<c\leq C<\infty$ such that for all $\theta\in\Delta$ and all $S\subset\{1,\ldots,K\}$, we have
@@
For any $\theta,\theta^{\prime}\in\Delta$, $y\in\{1,\ldots,K\}$, and $S\subset\{1,\ldots,K\}$, we have
@@
Consider $f(x)=\frac{1}{\alpha}(x^{\alpha}-1)-(x-1)$. We have $f(1)=0$ and $f^{\prime}(x)=x^{\alpha-1}-1\leq 0$ for $x\geq 1$. Hence we conclude. ∎
@@
Given $\epsilon>0$, there exists $m_{0}>0$ such that, for all $S\subset\{1,\ldots,K\}$ and $y\in\{1,\ldots,K\}$, $\ln h_{S,\epsilon}(y|\theta)$ is a concave function of $\vartheta$ that satisfies
@@
To cover both cases, take $m_{0}=\min\{\frac{\alpha}{6},\frac{1}{2}\}$. So the proof is complete. ∎
@@
$E_{\theta^{\ast}}[V_{t}(\theta,\theta^{\prime})]$ $=\sum_{S\subset\{1,\ldots,K\}}P_{\theta^{\ast}}(S_{t}=S)E_{\theta^{\ast}}[V_{t}(\theta,\theta^{\prime})|S_{t}=S]$
@@
Combining (37) and (38) and letting $c_{u}:=c\min_{S\subset\{1,\ldots,K\}}\lambda_{\min}(G_{S,\epsilon}^{\top}G_{S,\epsilon})$, we have
@@
$E_{\theta^{\ast}}\left(W_{t_{k}}|W_{t_{1}},\ldots,W_{t_{k-1}}\right)$ $=\sum_{S\subset\{1,\ldots,K\}}P_{\theta^{\ast}}(S_{t_{k}}=S|W_{t_{1}},\ldots,W_{t_{k-1}})E_{\theta^{\ast}}(W_{t_{k}}|S_{t_{k}}=S)$
$\leq\sum_{S\subset\{1,\ldots,K\}}P_{\theta^{\ast}}(S_{t_{k}}=S|W_{t_{1}},\ldots,W_{t_{k-1}})\left(1-c_{u}\right)$
@@
By Lemmas 8 and 9, we have $V_{t}(\theta,\theta^{\prime})\leq\|\theta-\theta^{\prime}\|^{2}$ and $E_{\theta^{\ast}}[V_{t}(\theta,\theta^{\prime})]\geq c_{u}\|\theta-\theta^{\prime}\|^{2}$. Thus we necessarily have $c_{u}<1$. Therefore, the recursion in (42) can be used until $k=1$ to obtain the desired result. ∎
@@
Next, we establish the $L$-smoothness of $h_{S,\epsilon}(y|\theta)$ as a function of $\theta$ for any $y\in\{1,\ldots,K\}$ and $S\subset\{1,\ldots,K\}$. Some technical lemmas are needed first.
@@
There exists an $L_{0}>0$ such that, for all $S\subset\{1,\ldots,K\}$ and $y\in\{1,\ldots,K\}$, the function $\ln h_{S,\epsilon}(y|\theta)$ is an $L_{0}$-smooth function of $\theta$. That is for all $\theta,\theta^{\prime}\in\Delta$, we have
@@
where the third line follows from Lemma 2. Hence we have arrived at the same inequality as (A.4.1). Hence, Lemma 14 holds with $L_{0}=\frac{1}{2c^{2}}$. ∎
@@
$E_{\theta^{\ast}}\left[\|\nabla_{\vartheta}\Phi_{n}(\theta^{\ast})\|^{2}\right]\leq\frac{1}{n}\max_{S\subset\{1,\ldots,K\}}\textup{Tr}\left[F(\theta^{\ast};S,\epsilon)\right].$
@@
$E_{\theta^{\ast}}\left[\nabla_{\vartheta}\ln h_{S_{t},\epsilon}(Y_{t}|\theta^{\ast})\right]$ $=\sum_{S\subset\{1,\ldots,K\}}P_{\theta^{\ast}}(S_{t}=S)E_{\theta^{\ast}}\left[\nabla_{\vartheta}\ln h_{S,\epsilon}(Y_{t}|\theta^{\ast})|S_{t}=S\right].$
@@
$\quad\quad=\sum_{S\subset\{1,\ldots,K\}}P_{\theta^{\ast}}(S_{t}=S)E_{\theta^{\ast}}\left[\nabla_{\vartheta}\ln h_{S,\epsilon}(Y_{t}|\theta^{\ast})\nabla_{\vartheta}\ln h_{S,\epsilon}(Y_{t}|\theta^{\ast})^{\top}|S_{t}=S\right]$
$\quad\quad=\sum_{S\subset\{1,\ldots,K\}}P_{\theta^{\ast}}(S_{t}=S)F(\theta^{\ast};S,\epsilon).$
@@
$E_{\theta^{\ast}}\left[\nabla_{\vartheta}\ln h_{S_{t},\epsilon}(Y_{t}|\theta^{\ast})|Y_{t^{\prime}},A_{t^{\prime}}\right]=\sum_{S\subset\{1,\ldots,K\}}P_{\theta^{\ast}}(S_{t}=S|Y_{t^{\prime}},S_{t^{\prime}})E_{\theta^{\ast}}\left[\nabla_{\vartheta}\ln h_{S,\epsilon}(Y_{t}|\theta^{\ast})|S_{t}=S\right]=0.$
@@
and hence we arrive at
$E_{\theta^{\ast}}\left[\nabla_{\vartheta}\Phi_{n}(\theta^{\ast})\nabla_{\vartheta}\Phi_{n}(\theta^{\ast})^{\top}\right]=\frac{1}{n^{2}}\sum_{t=1}^{n}\sum_{S\subset\{1,\ldots,K\}}P_{\theta^{\ast}}(S_{t}=S)\left[F(\theta^{\ast};S,\epsilon)\right]$ (48)
@@
$=\frac{1}{n^{2}}\sum_{t=1}^{n}\sum_{S\subset\{1,\ldots,K\}}P_{\theta^{\ast}}(S_{t}=S)\textup{Tr}(F(\theta^{\ast};S,\epsilon))$
$\leq\frac{1}{n}\max_{S\subset\{1,\ldots,K\}}\textup{Tr}\left(F(\theta^{\ast};S,\epsilon)\right),$

### v2

@@ 8 Conclusion
In this paper, we proposed a new adaptive framework, AdOBEst-LDP, for online estimation of the distribution of categorical data under the $\epsilon$-LDP constraint. AdOBEst-LDP, run with RRRR for randomization, encompasses both privatization of the sensitive data and accurate Bayesian estimation of population parameters from privatized data in a dynamic way. Our privatization mechanism (RRRR) is distinguished from the baseline approach (SRR) in a way that it operates on a smaller subset of the sample space rather than the entire sample space. We employed an adaptive approach to dynamically adjust the subset at each iteration, based on the knowledge about $\theta^{\ast}$ obtained from the past data. The selection of these subsets was guided by various alternative utility functions that we used throughout the paper. For the posterior sampling of $\theta$ at each iteration, we employed an efficient SGLD-based sampling scheme on a constrained region, namely the $K$-dimensional probability simplex. We distinguished this scheme from Gibbs sampling, which uses all of the historical data and is not scalable to large datasets.
@@
In the numerical experiments, we demonstrated that AdOBEst-LDP can estimate the population distribution more accurately than the non-adaptive approach under experimental settings with various privacy levels $\epsilon$ and degrees of evenness among the components of $\theta^{\ast}$. While the performance of AdOBEst-LDP is generally robust for all the utility functions considered in the paper, the utility function based on the probability of honest response can be preferred due to its much lower computational complexity than the other utility functions. Our experiments also showed that the accuracy of the adaptive approach is comparable to that of the semi-adaptive approach. However, the semi-adaptive approach requires adjusting its parameter $\alpha$ carefully, which makes it challenging to use.
@@
In a theoretical analysis, we showed that, regardless of whether the posterior sampling is conducted exactly or approximately, the posterior distribution targeted in AdOBEst-LDP converges to the true population parameter $\theta^{\ast}$. We also showed that, under exact posterior sampling, the best subset given utility function is selected with probability $1$ in the long run.
@@
It is important to note that the observations $\{Y_{t}\}_{t\geq 1}$ generated by AdOBEst-LDP are dependent. Therefore, the theoretical analysis presented in Section 6 can also be seen as a contribution to the literature on the convergence of posterior distributions with dependent data. Additionally, we have already highlighted an analogy between AdOBEst-LDP and Thompson sampling (Russo et al., , 2018). Both methods involve posterior sampling, and the subset selection step in AdOBEst-LDP can be viewed as analogous to the action selection step in reinforcement learning schemes. In this regard, we believe that the theoretical results may also inspire future research on the convergence of dynamic reinforcement learning algorithms, especially those based on Thompson sampling.
Categorical distributions serve as useful non-parametric discrete approximations of continuous distributions. As a potential future direction, AdOBEst-LDP could be adapted for non-parametric density estimation. A key challenge in this context would be determining how to partition the support domain of the data.
RRRR is a practical LDP mechanism with a subset parameter that adapts based on past data. It has been shown to outperform SRR when leveraging the knowledge of $\theta^{\ast}$. However, in this work, it is not proven that RRRR is the optimal $\epsilon$-LDP mechanism with respect to the utility functions considered. While the optimal $\epsilon$-LDP mechanism could be identified numerically by solving a constrained optimization problem—where the utility function is maximized under the LDP constraint—it may not have a closed-form solution for complex utility functions. A promising direction for future research would be to compare the optimal $\epsilon$-LDP mechanism with the $\epsilon$-LDP RRRR mechanism by analyzing their transition probability matrices and assessing the suboptimality of RRRR. Additionally, insights from the optimal $\epsilon$-LDP mechanism could inspire the development of new, tractable, and approximately optimal $\epsilon$-LDP mechanisms.
Acknowledgements
We thank our colleague Prof. Berrin Yanıkoğlu for reviewing the paper and providing insightful comments.
Appendix A Proofs
@@ A.1 Proofs for LDP of RRRR
for all possible $x,x^{\prime},y\in[K]$. When $S=\emptyset$, the proof is trivial; we focus on the non-trivial case $S\neq\emptyset$. For the non-trivial case, the transition probability $g_{S,\epsilon}(y|x)$ requires checking the ratio in (23) in $10$ different cases for $x,x^{\prime},y$ concerning their interrelation.
@@
    Since $\epsilon_{1}\leq\epsilon$, (23) holds.
@@
    Hence, we proved the right-hand side inequality. For the left-hand side, we have
@@
    Since $\epsilon_{1}\leq\epsilon$, (23) holds.
@@
    So (23) trivially holds.
We conclude the proof by noting that any other case left out is symmetric in $(x,x^{\prime})$ to one of the covered cases and, therefore, does not need to be checked separately. ∎
@@ A.2 Proofs about utility functions
$F(\theta;S,\epsilon)=\sum_{y=1}^{K}\frac{a_{y}^{\top}}{h_{S,\epsilon}(y|\theta)}\frac{a_{y}}{h_{S,\epsilon}(y|\theta)}h_{S,\epsilon}(y|\theta)=\sum_{y=1}^{K}a_{y}^{\top}\frac{1}{h_{S,\epsilon}(y|\theta)}a_{y}=A_{S,\epsilon}^{\top}D_{\theta}A_{S,\epsilon},$
@@
$G_{S,\epsilon}$ is invertible for all $S\subset[K]$ and $\epsilon>0$.
@@
The case $k=0$ is trivial since then $G_{S,\epsilon}=d_{1}I_{K}+d_{2}1_{K}1_{K}^{\top}$ is invertible. Hence, we focus on the case $0<k<K$. For this case, firstly, note that the matrices on the diagonal are invertible. So, by Weinstein–Aronszajn identity, for $G_{S,\epsilon}$ to be invertible, it suffices to show that the matrix
@@
Inside the square brackets is a scalar, therefore, $M$ in question is the sum of an identity matrix and a rank-$1$ matrix, which is invertible. Hence, $G_{S,\epsilon}$ is invertible. ∎
@@
where $c^{G}_{i}$ is the $i$’th column of $G_{S,\epsilon}$ for $i=1,\ldots,K$. Observe that $c^{A}_{i}$, $i=1,\ldots,K-1$ are linearly independent since any linear combination of those columns is in the form of
@@
Since the columns of $G_{S,\epsilon}$ are linearly independent, the linear combination above becomes 00 only if $a_{1}=\ldots=a_{K-1}=0$. This shows that the columns of $A_{S,\epsilon}$ are also linearly independent. Thus, we conclude that $A_{S,\epsilon}$ has rank $K-1$. Finally, since $D_{\theta}$ is diagonal with positive diagonal entries, $A_{S,\epsilon}^{\top}D_{\theta}A_{S,\epsilon}=A_{S,\epsilon}^{\top}D_{\theta}^{1/2}D_{\theta}^{1/2}A_{S,\epsilon}$ is positive definite, hence invertible. ∎
@@
The global maximization of $U_{6}$ over the set of all the subsets $S\subset[K]$ can be decomposed as
$\max_{S\subset[K]}U_{6}(\theta,S,\epsilon)=\max_{k\in\{0,\ldots,K-1\}}\left\{\max_{S\subset[K]:|S|=k}U_{6}(\theta,S,\epsilon)\right\}.$ (29)
@@ A.4.1 Preliminary results
Given $\epsilon\geq 0$, there exists constants $0<c\leq C<\infty$ such that for all $\theta\in\Delta$ and all $S\subset[K]$, we have
@@
For any $\theta,\theta^{\prime}\in\Delta$, $y\in[K]$, and $S\subset[K]$, we have
@@
Consider $f(x)=\frac{1}{\alpha}(x^{\alpha}-1)-(x-1)$. We have $f(1)=0$ and $f^{\prime}(x)=x^{\alpha-1}-1\leq 0$ for $x\geq 1$. Hence, we conclude. ∎
@@
Given $\epsilon>0$, there exists $m_{0}>0$ such that, for all $S\subset[K]$ and $y\in[K]$, $\ln h_{S,\epsilon}(y|\theta)$ is a concave function of $\vartheta$ that satisfies
@@
To cover both cases, take $m_{0}=\min\{\frac{\alpha}{6},\frac{1}{2}\}$. So, the proof is complete. ∎
@@
$E_{\theta^{\ast}}[V_{t}(\theta,\theta^{\prime})]$ $=\sum_{S\subset[K]}P_{\theta^{\ast}}(S_{t}=S)E_{\theta^{\ast}}[V_{t}(\theta,\theta^{\prime})|S_{t}=S]$
@@
Combining (38) and (39) and letting $c_{u}:=c\min_{S\subset[K]}\lambda_{\min}(G_{S,\epsilon}^{\top}G_{S,\epsilon})$, we have
@@
$E_{\theta^{\ast}}\left(W_{t_{k}}|W_{t_{1}},\ldots,W_{t_{k-1}}\right)$ $=\sum_{S\subset[K]}P_{\theta^{\ast}}(S_{t_{k}}=S|W_{t_{1}},\ldots,W_{t_{k-1}})E_{\theta^{\ast}}(W_{t_{k}}|S_{t_{k}}=S)$
$\leq\sum_{S\subset[K]}P_{\theta^{\ast}}(S_{t_{k}}=S|W_{t_{1}},\ldots,W_{t_{k-1}})\left(1-c_{u}\right)$
@@
By Lemmas 8 and 9, we have $V_{t}(\theta,\theta^{\prime})\leq\|\theta-\theta^{\prime}\|^{2}$ and $E_{\theta^{\ast}}[V_{t}(\theta,\theta^{\prime})]\geq c_{u}\|\theta-\theta^{\prime}\|^{2}$. Thus, we necessarily have $c_{u}<1$. Therefore, the recursion in (43) can be used until $k=1$ to obtain the desired result. ∎
@@
Next, we establish the $L$-smoothness of $h_{S,\epsilon}(y|\theta)$ as a function of $\theta$ for any $y\in[K]$ and $S\subset[K]$. Some technical lemmas are needed first.
@@
There exists an $L_{0}>0$ such that, for all $S\subset[K]$ and $y\in[K]$, the function $\ln h_{S,\epsilon}(y|\theta)$ is an $L_{0}$-smooth function of $\theta$. That is for all $\theta,\theta^{\prime}\in\Delta$, we have
@@
where the third line follows from Lemma 2. Hence, we have arrived at the same inequality as (A.4.1). Hence, Lemma 14 holds with $L_{0}=\frac{1}{2c^{2}}$. ∎
@@
$E_{\theta^{\ast}}\left[\|\nabla_{\vartheta}\Phi_{n}(\theta^{\ast})\|^{2}\right]\leq\frac{1}{n}\max_{S\subset[K]}\textup{Tr}\left[F(\theta^{\ast};S,\epsilon)\right].$
@@
$E_{\theta^{\ast}}\left[\nabla_{\vartheta}\ln h_{S_{t},\epsilon}(Y_{t}|\theta^{\ast})\right]$ $=\sum_{S\subset[K]}P_{\theta^{\ast}}(S_{t}=S)E_{\theta^{\ast}}\left[\nabla_{\vartheta}\ln h_{S,\epsilon}(Y_{t}|\theta^{\ast})|S_{t}=S\right].$
@@
$\quad\quad=\sum_{S\subset[K]}P_{\theta^{\ast}}(S_{t}=S)E_{\theta^{\ast}}\left[\nabla_{\vartheta}\ln h_{S,\epsilon}(Y_{t}|\theta^{\ast})\nabla_{\vartheta}\ln h_{S,\epsilon}(Y_{t}|\theta^{\ast})^{\top}|S_{t}=S\right]$
$\quad\quad=\sum_{S\subset[K]}P_{\theta^{\ast}}(S_{t}=S)F(\theta^{\ast};S,\epsilon).$
@@
$E_{\theta^{\ast}}\left[\nabla_{\vartheta}\ln h_{S_{t},\epsilon}(Y_{t}|\theta^{\ast})|Y_{t^{\prime}},A_{t^{\prime}}\right]=\sum_{S\subset[K]}P_{\theta^{\ast}}(S_{t}=S|Y_{t^{\prime}},S_{t^{\prime}})E_{\theta^{\ast}}\left[\nabla_{\vartheta}\ln h_{S,\epsilon}(Y_{t}|\theta^{\ast})|S_{t}=S\right]=0.$
@@
and hence, we arrive at
$E_{\theta^{\ast}}\left[\nabla_{\vartheta}\Phi_{n}(\theta^{\ast})\nabla_{\vartheta}\Phi_{n}(\theta^{\ast})^{\top}\right]=\frac{1}{n^{2}}\sum_{t=1}^{n}\sum_{S\subset[K]}P_{\theta^{\ast}}(S_{t}=S)\left[F(\theta^{\ast};S,\epsilon)\right]$ (49)
@@
$=\frac{1}{n^{2}}\sum_{t=1}^{n}\sum_{S\subset[K]}P_{\theta^{\ast}}(S_{t}=S)\textup{Tr}(F(\theta^{\ast};S,\epsilon))$
$\leq\frac{1}{n}\max_{S\subset[K]}\textup{Tr}\left(F(\theta^{\ast};S,\epsilon)\right),$